OpenStudy (anonymous):

Suppose that y varies directly with x. Find y when x=5 if y=-6 when x=2

OpenStudy (anonymous):

y=-6 when x=2:use ratio and proportion to this -6/2=y/5 by cross multiplying, y=5(-3) y=-15